# Center-freeness of finite-step solvable groups arising from anabelian geometry

### v1

@@ 1.2. A first computation of a centralizer in a free $\Sigma pro$duct
For a profinite group $G$ and an element $g\in G$, we define
$\mathrm{C}_{G}(g)\coloneq\{h\in G\mid hgh^{-1}=g\},$
and call it the centralizer of $g$ in $G$. (Note that this group is closed in $G$, and hence profinite.) A slightly different version of the following proposition first appeared in [11, Lemma 2.1.2], where it was used to prove the center-freeness of free discrete groups. We generalize it to our setting as follows:
@@
$\overline{(\,\cdot\,)}:\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\rightarrow\mathrm{M}_{u}(\mathbb{Z}/\ell\mathbb{Z})$
be the reduction morphism induced by $\mathbb{Z}/\ell^{\sigma}\mathbb{Z}\twoheadrightarrow\mathbb{Z}/\ell\mathbb{Z}$. If $E\in\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$ satisfies $\tilde{n}E=0$, then $\overline{E}=0$ in $\mathrm{M}_{u}(\mathbb{Z}/\ell\mathbb{Z})$.
@@
On the other hand, $\ell\,\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$ is exactly the kernel of the reduction morphism $\overline{(\cdot)}$. Thus $\overline{E}=0$. ∎
@@
Let $\Omega=\mathcal{C}*P$ be the free $\Sigma pro$duct (see [12, Proposition 9.1.2]) of a procyclic pro-$\Sigma$ group $\mathcal{C}$, topologically generated by an element $x$, and a pro-$\Sigma$ group $P$. Let $m\in\mathbb{Z}_{\geq 2}$. Then, for any $n\in\mathbb{Z}$ such that $x^{n}\neq 1$ in $\mathcal{C}$, one has
$\mathrm{C}_{\Omega^{m}}(x^{n})\subset\langle x\rangle\cdot(\Omega^{m})^{[m-1]}$ (1.1)
as a subgroup of $\Omega^{m}$, where $\langle x\rangle$ denotes the closed subgroup of $\Omega^{m}$ generated by the image of $x$.
@@
Since $x^{-1}$ is also a topological generator of $\mathcal{C}$, we may assume that $n\geq 1$. To prove 1.1, it suffices to show that
$\rho\left(\mathrm{C}_{\Omega^{m}}(x^{n})\right)\subset\langle\rho(x)\rangle$ (1.2)
for any continuous surjection $\rho:\Omega^{m}\twoheadrightarrow G$ onto a finite group $G$ that factors through the natural projection $\Omega^{m}\twoheadrightarrow\Omega^{m-1}$. Since $\Omega=\mathcal{C}*P$, we have $\Omega^{\mathrm{ab}}\cong\mathcal{C}^{\mathrm{ab}}\times P^{\mathrm{ab}}\cong\mathcal{C}\times P^{\mathrm{ab}}$. In particular, the composition of the natural morphisms $\mathcal{C}\rightarrow\Omega\rightarrow\Omega^{m}$ is injective, and the family of surjections $\rho$ such that $\rho(x^{n})\neq 1$ is cofinal. Therefore, we may assume that $\rho(x^{n})\neq 1$.
To prove 1.2, it suffices to construct a profinite group $\tilde{G}$ and a factorization
@@
$\phi\bigl(\mathrm{C}_{\tilde{G}}(\psi(x)^{n})\bigr)\subset\bigl\langle\phi\circ\psi(x)\bigr\rangle.$ (1.3)
@@
$\rho\bigl(\mathrm{C}_{\Omega^{m}}(x^{n})\bigr)=(\phi\circ\psi)\bigl(\mathrm{C}_{\Omega^{m}}(x^{n})\bigr)\subset\phi\bigl(\mathrm{C}_{\tilde{G}}(\psi(x)^{n})\bigr)\subset\bigl\langle\phi\circ\psi(x)\bigr\rangle=\langle\rho(x)\rangle.$
@@
$\tilde{G}\coloneq\left\{\left(\begin{array}[]{cc}A&B\\ 0&C\end{array}\right)\in\operatorname{GL}_{2u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\,\middle|\,A\in G,\ B\in\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z}),\ C\in\langle\rho(x)\rangle\right\}.$
By construction, $\tilde{G}$ fits into the short exact sequence
$1\to(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})^{\oplus u^{2}}\to\tilde{G}\to G\times\langle\rho(x)\rangle\to 1.$
Since $\rho:\Omega^{m}\twoheadrightarrow G$ factors through $\Omega^{m}\twoheadrightarrow\Omega^{m-1}$, the group $G\times\langle\rho(x)\rangle$ is (m1)(m-1)-step solvable. Therefore, $\tilde{G}$ is an $m$-step solvable pro-$\Sigma$ group. The surjection $\Omega\to\Omega^{m}\xrightarrow{\rho}G$ extends to a morphism $\psi:\Omega\to\tilde{G}$, defined by
$x\mapsto\left(\begin{array}[]{cc}\rho(x)&\rho(x)\\ 0&\rho(x)\end{array}\right),\qquad p\mapsto\left(\begin{array}[]{cc}\rho(p)&0\\ 0&I_{u}\end{array}\right)\qquad\text{for all }p\in P.$
@@
Finally, we show the desired property 1.3. Let $y\in\mathrm{C}_{\tilde{G}}(\psi(x)^{n})$ and write
@@
are the same. By comparing the top-right blocks, we obtain sn(AC)=0sn(A-C)=0 in $\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$. By Lemma 1.2, this implies $\overline{A}=\overline{C}$ in $\mathrm{M}_{u}(\mathbb{Z}/\ell\mathbb{Z})$. Since $G\hookrightarrow\operatorname{GL}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\twoheadrightarrow\operatorname{GL}_{u}(\mathbb{Z}/\ell\mathbb{Z})$ is still injective, we conclude that $\phi(y)=A=C$ in $G$. Therefore, $\phi(y)\in\langle\rho(x)\rangle$, i.e., 1.3 holds. This completes the proof. ∎
@@ 1.3. Proof of the slimness of free $m$-step solvable groups
Using the above ingredients, we compute explicitly the centralizer of a free generator of a (possibly infinitely generated) free $m$-step solvable pro-$\Sigma$ group, and obtain the slimness of such profinite groups.
@@
Let $\mathcal{F}$ be a free pro-$\Sigma$ group of finite rank $r$ with free generating set $X$. Then for any non-zero integer $n\in\mathbb{Z}_{\neq 0}$ and any $x\in X$, the element $\overline{x}^{n}-1$ is a non-zero-divisor in $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$, where $\overline{x}$ is the image of $x$ in $\mathcal{F}^{\mathrm{ab}}$.
@@
Denote by $\mathbb{Z}(\Sigma)_{\geq 1}$ the set of all positive integers whose prime factors lie in $\Sigma$. We may assume that $n\geq 1$ as $\overline{x}^{-n}-1=-\overline{x}^{-n}(\overline{x}^{\,n}-1)$ in $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$. Let us show that if $y\in\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$ satisfies $(\overline{x}^{n}-1)y=0$, then $y=0$.
@@
where $H\cong\mathbb{Z}_{\Sigma}^{r-1}$ is the free abelian factor generated by the images of $X\setminus\{x\}$, and the factor $\mathbb{Z}_{\Sigma}$ corresponds to $\overline{x}$. Put
@@
For each $N\in\mathbb{Z}(\Sigma)_{\geq 1}$, let $C_{N}\coloneq\langle\overline{x}_{N}\mid(\overline{x}_{N})^{N}=1\rangle\cong\mathbb{Z}/N\mathbb{Z}$. Then, by the definition of the completed group algebra and the above decomposition, we have
@@
Here, we may regard $\overline{x}$ as the projective limit of $(\overline{x}_{N})_{N}$.
Write $y_{N}$ for the image of $y$ in $A[C_{N}]$. Since $\{1,\overline{x}_{N},\cdots,(\overline{x}_{N})^{N-1}\}$ is an AA-basis of $A[C_{N}]$, there exists a unique $c_{i}^{(N)}\in A$ such that
$y_{N}=\sum_{i=0}^{N-1}c_{i}^{(N)}\overline{x}_{N}^{i}.$
The equation $(\overline{x}^{n}-1)y=0$ implies $((\overline{x}_{N})^{n}-1)y_{N}=0$ for all NN, and hence
$0=((\overline{x}_{N})^{n}-1)\left(\sum_{i=0}^{N-1}c_{i}^{(N)}\overline{x}_{N}^{i}\right)=\sum_{i=0}^{N-1}\bigl(c_{i-n}^{(N)}-c_{i}^{(N)}\bigr)\overline{x}_{N}^{i},$
where indices $\{i\}$ of $c_{i}^{(N)}$ are taken in $\mathbb{Z}/N\mathbb{Z}$. By AA-linear independence of $\{\overline{x}_{N}^{i}\}$, we obtain $c_{i-n}^{(N)}=c_{i}^{(N)}$ for all $i\in\mathbb{Z}/N\mathbb{Z}$. In other words, $c_{i}^{(N)}$ is constant on cosets of the subgroup $\langle n\rangle\subset\mathbb{Z}/N\mathbb{Z}$.
Let $n=n_{\Sigma}\cdot n_{\Sigma^{\prime}}$ be the unique decomposition such that $n_{\Sigma}\in\mathbb{Z}(\Sigma)_{\geq 1}$ and that $n_{\Sigma^{\prime}}$ is coprime to all primes in $\Sigma$. Fix $M\in\mathbb{Z}(\Sigma)_{\geq 1}$ such that $n_{\Sigma}\mid M$, and let $k\in\mathbb{Z}(\Sigma)_{\geq 1}$ be arbitrary. As $n_{\Sigma^{\prime}}$ and kMkM are coprime to each other, we have $\langle n\rangle=\langle n_{\Sigma}\rangle\subset\mathbb{Z}/kM\mathbb{Z}$. Hence we may apply the above result for NN to kMkM, which gives $c_{i-n_{\Sigma}}^{(kM)}=c_{i}^{(kM)}$ for all $i\in\mathbb{Z}/kM\mathbb{Z}$. Therefore, by $n_{\Sigma}\mid M$, we obtain
@@
for all $i\in\mathbb{Z}/kM\mathbb{Z}$. Let $\pi:A[C_{kM}]\to A[C_{M}],\ \overline{x}_{kM}\mapsto\overline{x}_{M}$, be the natural projection induced by $\mathbb{Z}/kM\mathbb{Z}\twoheadrightarrow\mathbb{Z}/M\mathbb{Z}$. By $\pi(\overline{x}_{kM})=\overline{x}_{M}$ and 1.4, we have
$y_{M}=\pi(y_{kM})=\sum_{i=0}^{kM-1}c_{i}^{(kM)}\,\pi\bigl(\overline{x}_{kM}^{i}\bigr)=\sum_{i=0}^{M-1}\Bigl(\sum_{j=0}^{k-1}c_{i+jM}^{(kM)}\Bigr)\overline{x}_{M}^{i}=\sum_{i=0}^{M-1}\bigl(k\cdot c_{i}^{(kM)}\bigr)\overline{x}_{M}^{i}.$
Comparing this with $y_{M}=\sum_{i=0}^{M-1}c_{i}^{(M)}\overline{x}_{M}^{i}$, we obtain
@@
for all $i\in\mathbb{Z}/M\mathbb{Z}$. By running over all $k\in\mathbb{Z}(\Sigma)_{\geq 1}$ and using the fact $\bigcap_{k}kA=\{0\}$, we obtain $y_{M}=0$. Since the set $\{M\in\mathbb{Z}(\Sigma)_{\geq 1}\mid n_{\Sigma}\mid M\}$ is cofinal in $\mathbb{Z}(\Sigma)_{\geq 1}$, it follows that $y=0$. This completes the proof. ∎
@@
Let $\mathcal{F}$ be a (possibly infinitely generated) free pro-$\Sigma$ group of rank $r$ with free generating set $X$. Let $m\in\mathbb{Z}_{\geq 2}$. Then for any non-zero integer $n\in\mathbb{Z}_{\neq 0}$ and any $x\in X$, one has
$\mathrm{C}_{\mathcal{F}^{m}}(x^{n})=\langle x\rangle.$
@@
If $r=1$, the assertion is clear. Hence we may assume $r\neq 1$. Fix $x\in X$. We divide the proof into three cases: the case $m=2$ with $r$ finite; the case of general $m$ with $r$ finite; and the general case.
First, we assume that $m=2$ and $r$ is finite. By Proposition 1.3 and the fact that $\langle x\rangle\subset\mathrm{C}_{\mathcal{F}^{2}}(x^{n})$, we obtain $\mathrm{C}_{\mathcal{F}^{2}}(x^{n})=\langle x\rangle\cdot(\mathrm{C}_{\mathcal{F}^{2}}(x^{n})\cap(\mathcal{F}^{2})^{[1]})$. Therefore, it suffices to show that
$\mathrm{C}_{\mathcal{F}^{2}}(x^{n})\cap(\mathcal{F}^{2})^{[1]}=1.$ (1.5)
@@
$\mathrm{C}_{\mathcal{F}^{2}}(x^{n})\cap(\mathcal{F}^{2})^{[1]}$ $=$ $\{u\in(\mathcal{F}^{2})^{[1]}\mid x^{n}ux^{-n}=u\}$
$=$ $\ker\bigl((\overline{x}^{n}-1):(\mathcal{F}^{2})^{[1]}\to(\mathcal{F}^{2})^{[1]}\bigr)$
$\subset$ $\ker\bigl((\overline{x}^{n}-1):\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]^{\oplus r}\to\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]^{\oplus r}\bigr),$
where $\overline{x}$ is the image of $x$ in $\mathcal{F}^{\mathrm{ab}}$. By Lemma 1.4, the element $\overline{x}^{n}-1$ is a non-zero-divisor in $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$, and hence multiplication by $\overline{x}^{n}-1$ is injective. Therefore, the last kernel is trivial, and hence the equation 1.5 follows. This proves
$\mathrm{C}_{\mathcal{F}^{2}}(x^{n})=\langle x\rangle$
@@
Next, assume that $r$ is finite and proceed by induction on $m\in\mathbb{Z}_{\geq 2}$. The case of $m=2$ is already proved. Suppose that $m>2$ and that the assertion holds for m1m-1. As in the case $m=2$, by Proposition 1.3, it suffices to show that
$\mathrm{C}_{\mathcal{F}^{m}}(x^{n})\cap(\mathcal{F}^{m})^{[m-1]}=1.$ (1.6)
Let $g$ be an element of the left-hand side of 1.6. Let $H$ be an open normal subgroup of $\mathcal{F}^{m}$ containing $(\mathcal{F}^{m})^{[1]}$. Since

### v3

@@ 1.2. A computation of a centralizer in a free $\Sigma pro$duct
A slightly different version of the following proposition first appeared in [11, Lemma 2.1.2], where it was used to prove the center-freeness of free discrete groups. We generalize it to our setting as follows:
@@
$\Phi\colon\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\to\mathrm{M}_{u}(\mathbb{Z}/\ell\mathbb{Z})$
be the reduction morphism induced by $\mathbb{Z}/\ell^{\sigma}\mathbb{Z}\twoheadrightarrow\mathbb{Z}/\ell\mathbb{Z}$. If $E\in\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$ satisfies $\tilde{n}E=0$, then $\Phi(E)=0$.
@@
On the other hand, the subgroup $\ell\,\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$ is exactly the kernel of $\Phi$. Thus $\Phi(E)=0$. ∎
@@
Let $\Omega=\mathcal{C}*P$ be the free $\Sigma pro$duct (see [12, Proposition 9.1.2]) of a procyclic pro-$\Sigma$ group $\mathcal{C}$, topologically generated by an element $x$, and a pro-$\Sigma$ group $P$. Let $m\in\mathbb{Z}_{\geq 2}$. Then, for any $n\in\mathbb{Z}$ such that $x^{n}\neq 1$ in $\mathcal{C}$, we have
$\mathrm{C}_{\Omega^{m}}\left(x^{n}\right)\subset\overline{\left\langle x\right\rangle}\cdot(\Omega^{m})^{[m-1]}$ (1.1)
as a subgroup of $\Omega^{m}$, where $\overline{\left\langle x\right\rangle}$ denotes the closed subgroup of $\Omega^{m}$ topologically generated by the image of $x$.
@@
Since $x^{-1}$ is also a topological generator of $\mathcal{C}$, we may assume that $n\geq 1$. To prove (1.1), it suffices to show that, for any continuous surjection $\rho:\Omega^{m}\twoheadrightarrow G$ onto a finite group $G$ that factors through the natural projection $\Omega^{m}\twoheadrightarrow\Omega^{m-1}$, we have
$\rho\left(\mathrm{C}_{\Omega^{m}}\left(x^{n}\right)\right)\subset\overline{\left\langle\rho(x)\right\rangle}.$ (1.2)
Since $\Omega=\mathcal{C}*P$, we have $\Omega^{\mathrm{ab}}\cong\mathcal{C}^{\mathrm{ab}}\times P^{\mathrm{ab}}\cong\mathcal{C}\times P^{\mathrm{ab}}$. In particular, the composition of the natural morphisms $\mathcal{C}\to\Omega\to\Omega^{m}$ is injective, and the family of surjections $\rho$ such that $\rho(x^{n})\neq 1$ is cofinal. Therefore, we may assume that $\rho(x^{n})\neq 1$ in the above.
To prove (1.2), it suffices to construct a profinite group $\tilde{G}$ and a factorization
@@
$\phi\bigl(\mathrm{C}_{\tilde{G}}\left(\psi(x)^{n}\right)\bigr)\subset\overline{\left\langle\phi\circ\psi(x)\right\rangle}.$ (1.3)
@@
$\rho\bigl(\mathrm{C}_{\Omega^{m}}\left(x^{n}\right)\bigr)=(\phi\circ\psi)\bigl(\mathrm{C}_{\Omega^{m}}\left(x^{n}\right)\bigr)\subset\phi\bigl(\mathrm{C}_{\tilde{G}}\left(\psi(x)^{n}\right)\bigr)\subset\overline{\left\langle\phi\circ\psi(x)\right\rangle}=\overline{\left\langle\rho(x)\right\rangle}.$
@@
$\tilde{G}\coloneq\left\{\left(\begin{array}[]{cc}A&B\\ 0&C\end{array}\right)\in\operatorname{GL}_{2u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\,\middle|\,A\in G,\ B\in\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z}),\ C\in\overline{\left\langle\rho(x)\right\rangle}\right\}.$
By construction, the group $\tilde{G}$ fits into the short exact sequence
$1\to(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})^{\oplus u^{2}}\to\tilde{G}\to G\times\overline{\left\langle\rho(x)\right\rangle}\to 1.$
Since $\rho:\Omega^{m}\twoheadrightarrow G$ factors through $\Omega^{m}\twoheadrightarrow\Omega^{m-1}$, the group $G\times\overline{\left\langle\rho(x)\right\rangle}$ is (m1)(m-1)-step solvable. Therefore, the group $\tilde{G}$ is an $m$-step solvable pro-$\Sigma$ group. The surjection $\Omega\to\Omega^{m}\xrightarrow{\rho}G$ extends to a morphism $\psi:\Omega\to\tilde{G}$, defined by
$x\mapsto\left(\begin{array}[]{cc}\rho(x)&\rho(x)\\ 0&\rho(x)\end{array}\right),\qquad p\mapsto\left(\begin{array}[]{cc}\rho(p)&0\\ 0&I_{u}\end{array}\right)\qquad\text{for each }p\in P.$
@@
Finally, we show the desired property (1.3). Let $y\in\mathrm{C}_{\tilde{G}}\left(\psi(x)^{n}\right)$ and write
@@
coincide. By comparing the top-right blocks, we obtain sn(AC)=0sn(A-C)=0 in $\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$. By Lemma 1.2, this implies $\Phi(A)=\Phi(C)$ in $\mathrm{M}_{u}(\mathbb{Z}/\ell\mathbb{Z})$. Since $G\hookrightarrow\operatorname{GL}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\twoheadrightarrow\operatorname{GL}_{u}(\mathbb{Z}/\ell\mathbb{Z})$ is still injective, we conclude that $\phi(y)=A=C$ in $G$. Therefore, we have $\phi(y)\in\overline{\left\langle\rho(x)\right\rangle}$. Thus (1.3) holds. This completes the proof. ∎
@@ 1.3. Proof of the slimness of free $m$-step solvable groups
Using the above ingredients, we compute explicitly the centralizer of a free generator in a (possibly infinitely generated) free $m$-step solvable pro-$\Sigma$ group and deduce the slimness of such profinite groups.
@@
Let $\mathcal{F}$ be a free pro-$\Sigma$ group of finite rank $r$ with free generating set $X$. For any nonzero integer $n\in\mathbb{Z}$ and any $x\in X$, the element $\underline{x}^{n}-1$ is a nonzero divisor in $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$, where $\underline{x}$ is the image of $x$ in $\mathcal{F}^{\mathrm{ab}}$.
@@
Denote by $\mathbb{Z}(\Sigma)_{\geq 1}$ the set of all positive integers whose prime factors lie in $\Sigma$. We may assume that $n\geq 1$ since $\underline{x}^{-n}-1=-\underline{x}^{-n}(\underline{x}^{\,n}-1)$ in $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$. We show that if $y\in\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$ satisfies $(\underline{x}^{n}-1)y=0$, then $y=0$.
@@
where $H\cong\mathbb{Z}_{\Sigma}^{r-1}$ is the free abelian factor generated by the images of $X\setminus\{x\}$, and the factor $\mathbb{Z}_{\Sigma}$ corresponds to $\underline{x}$. Put
@@
For each $N\in\mathbb{Z}(\Sigma)_{\geq 1}$, let $C_{N}\coloneq\langle\underline{x}_{N}\mid(\underline{x}_{N})^{N}=1\rangle\cong\mathbb{Z}/N\mathbb{Z}$. Then, by the definition of the completed group algebra and the above decomposition, we have
@@
Here, we may regard $\underline{x}$ as the projective limit of $(\underline{x}_{N})_{N}$.
Write $y_{N}$ for the image of $y$ in $A[C_{N}]$. Since $\{1,\underline{x}_{N},\cdots,(\underline{x}_{N})^{N-1}\}$ is an $A$-basis of $A[C_{N}]$, there exists a unique $c_{i}^{(N)}\in A$ such that
$y_{N}=\sum_{i=0}^{N-1}c_{i}^{(N)}\underline{x}_{N}^{i}.$
The equation $(\underline{x}^{n}-1)y=0$ implies $((\underline{x}_{N})^{n}-1)y_{N}=0$ for any $N$, and hence
$0=((\underline{x}_{N})^{n}-1)\left(\sum_{i=0}^{N-1}c_{i}^{(N)}\underline{x}_{N}^{i}\right)=\sum_{i=0}^{N-1}\bigl(c_{i-n}^{(N)}-c_{i}^{(N)}\bigr)\underline{x}_{N}^{i},$
where the indices $i$ of $c_{i}^{(N)}$ are taken in $\mathbb{Z}/N\mathbb{Z}$. By $A$-linear independence of $\{\underline{x}_{N}^{i}\}$, we obtain $c_{i-n}^{(N)}=c_{i}^{(N)}$ for each $i\in\mathbb{Z}/N\mathbb{Z}$. In other words, the coefficients $c_{i}^{(N)}$ are constant on cosets of the subgroup $\langle n\rangle\subset\mathbb{Z}/N\mathbb{Z}$.
Let $n=n_{\Sigma}\cdot n_{\Sigma^{\prime}}$ be the unique decomposition such that $n_{\Sigma}\in\mathbb{Z}(\Sigma)_{\geq 1}$ and that $n_{\Sigma^{\prime}}$ is coprime to all primes in $\Sigma$. Fix $M\in\mathbb{Z}(\Sigma)_{\geq 1}$ such that $n_{\Sigma}\mid M$, and let $k\in\mathbb{Z}(\Sigma)_{\geq 1}$ be arbitrary. As $n_{\Sigma^{\prime}}$ and $kM$ are coprime to each other, we have $\langle n\rangle=\langle n_{\Sigma}\rangle\subset\mathbb{Z}/kM\mathbb{Z}$. Hence we may apply the above result with $N=kM$, which gives $c_{i-n_{\Sigma}}^{(kM)}=c_{i}^{(kM)}$ for each $i\in\mathbb{Z}/kM\mathbb{Z}$. Therefore, by $n_{\Sigma}\mid M$, we obtain
@@
for each $i\in\mathbb{Z}/kM\mathbb{Z}$. Let $\pi:A[C_{kM}]\to A[C_{M}],\ \underline{x}_{kM}\mapsto\underline{x}_{M}$, be the natural projection induced by $\mathbb{Z}/kM\mathbb{Z}\twoheadrightarrow\mathbb{Z}/M\mathbb{Z}$. By $\pi(\underline{x}_{kM})=\underline{x}_{M}$ and (1.4), we have
$y_{M}=\pi(y_{kM})=\sum_{i=0}^{kM-1}c_{i}^{(kM)}\,\pi\bigl(\underline{x}_{kM}^{i}\bigr)=\sum_{i=0}^{M-1}\Bigl(\sum_{j=0}^{k-1}c_{i+jM}^{(kM)}\Bigr)\underline{x}_{M}^{i}=\sum_{i=0}^{M-1}\bigl(k\cdot c_{i}^{(kM)}\bigr)\underline{x}_{M}^{i}.$
Comparing this with $y_{M}=\sum_{i=0}^{M-1}c_{i}^{(M)}\underline{x}_{M}^{i}$, we obtain
@@
for each $i\in\mathbb{Z}/M\mathbb{Z}$. By running over all $k\in\mathbb{Z}(\Sigma)_{\geq 1}$ and using the fact $\bigcap_{k}kA=\{0\}$, we obtain $y_{M}=0$. Since the set $\{M\in\mathbb{Z}(\Sigma)_{\geq 1}\mid n_{\Sigma}\mid M\}$ is cofinal in $\mathbb{Z}(\Sigma)_{\geq 1}$, it follows that $y=0$. This completes the proof. ∎
@@
Let $\mathcal{F}$ be a (possibly infinitely generated) free pro-$\Sigma$ group of rank $r$ with free generating set $X$. Let $m\in\mathbb{Z}_{\geq 2}$. Then, for any nonzero integer $n\in\mathbb{Z}$ and any $x\in X$, we have
$\mathrm{C}_{\mathcal{F}^{m}}\left(x^{n}\right)=\overline{\left\langle x\right\rangle}.$
@@
If $r=1$, the assertion is clear. Hence we may assume $r\neq 1$. Fix $x\in X$. We divide the proof into three cases: $m=2$ with finite $r$; general $m$ with finite $r$; and the case $r=\infty$.
First, we assume that $m=2$ and $r$ is finite. By Proposition 1.3 and the fact that $\overline{\left\langle x\right\rangle}\subset\mathrm{C}_{\mathcal{F}^{2}}\left(x^{n}\right)$, we obtain $\mathrm{C}_{\mathcal{F}^{2}}\left(x^{n}\right)=\overline{\left\langle x\right\rangle}\cdot(\mathrm{C}_{\mathcal{F}^{2}}\left(x^{n}\right)\cap(\mathcal{F}^{2})^{[1]})$. Therefore, it suffices to show that
$\mathrm{C}_{\mathcal{F}^{2}}\left(x^{n}\right)\cap(\mathcal{F}^{2})^{[1]}=1.$ (1.5)
@@
$\mathrm{C}_{\mathcal{F}^{2}}\left(x^{n}\right)\cap(\mathcal{F}^{2})^{[1]}$ $=$ $\{u\in(\mathcal{F}^{2})^{[1]}\mid x^{n}ux^{-n}=u\}$
$=$ $\ker\bigl((\underline{x}^{n}-1):(\mathcal{F}^{2})^{[1]}\to(\mathcal{F}^{2})^{[1]}\bigr)$
$\subset$ $\ker\bigl((\underline{x}^{n}-1):\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]^{\oplus r}\to\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]^{\oplus r}\bigr),$
where $\underline{x}$ is the image of $x$ in $\mathcal{F}^{\mathrm{ab}}$. By Lemma 1.4, the element $\underline{x}^{n}-1$ is a nonzero divisor in $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$, and hence multiplication by $\underline{x}^{n}-1$ is injective. Therefore, the last kernel is trivial, and hence the equation (1.5) follows. This proves
$\mathrm{C}_{\mathcal{F}^{2}}\left(x^{n}\right)=\overline{\left\langle x\right\rangle}$
@@
Next, assume that $r$ is finite and proceed by induction on $m\in\mathbb{Z}_{\geq 2}$. The case of $m=2$ is already proved. Assume that $m>2$ and that the assertion holds for m1m-1. As in the case $m=2$, by Proposition 1.3, it suffices to show that
$\mathrm{C}_{\mathcal{F}^{m}}\left(x^{n}\right)\cap(\mathcal{F}^{m})^{[m-1]}=1.$ (1.6)
Let $g$ be an element of the left-hand side of (1.6). Let $H$ be an open normal subgroup of $\mathcal{F}^{m}$ that contains $(\mathcal{F}^{m})^{[1]}$. Since